# Initialization is Critical to Whether Transformers Fit Composite Functions by Inference or Memorizing

### v2

@@ Abstract
Transformers have shown impressive capabilities across various tasks, but their performance on compositional problems remains a topic of debate. In this work, we investigate the mechanisms of how transformers behave on unseen compositional tasks. We discover that the parameter initialization scale plays a critical role in determining whether the model learns inferential solutions, which capture the underlying compositional primitives, or symmetric solutions, which simply memorize mappings without understanding the compositional structure. By analyzing the information flow and vector representations within the model, we reveal the distinct mechanisms underlying these solution types. We further find that inferential solutions exhibit low complexity bias, which we hypothesize is a key factor enabling them to learn individual mappings for single anchors. Building upon the understanding of these mechanisms, we can predict the learning behavior of models with different initialization scales when faced with data of varying complexity. Our findings provide valuable insights into the role of initialization scale in shaping the type of solution learned by transformers and their ability to learn and generalize compositional tasks.
@@ 1 Introduction
Large-scale transformers (Vaswani et al., 2017) demonstrate unprecedented capabilities (Achiam et al., 2023; Brown et al., 2020; Choi et al., 2021; Teubner et al., 2023), even noted as "sparks of AGI" (Bubeck et al., 2023). However, their performance on compositional tasks, which are considered a key property of human cognition (Marcus, 2003; Smolensky et al., 2022), has long been the subject of intense debate, especially in terms of the out-of-distribution (OOD) generalization ability. This raises critical open questions about how to faithfully interpret transformers’ capabilities on compositional tasks. Can transformers learn the underlying compositional primitives within the data, or do they merely learn input-output mappings? When transformers produce incorrect outputs on compositional tasks, do their responses follow any discernible patterns or logic?
@@
Building upon our understanding of these mechanisms, we predict and experimentally validate that models with small initialization scales require more training steps to fit data with larger inferential complexity, defined as the number of anchor pair groups that do not satisfy the inferential mapping, while models with large initialization scales use similar training steps regardless of inferential complexity.
Our work highlights the crucial role of initialization scale in shaping the solutions learned by transformers on compositional tasks, enabling them to better capture compositional structures and generalize to unseen tasks. Furthermore, adapting the initialization scale to the specific task type could be promising: larger scales for memorization tasks such as text memorization, and smaller scales for inference tasks such as code generation. Unless otherwise specified, our main text primarily focuses on a single-head attention model to facilitate a clearer understanding of the underlying mechanisms. We further extend our experiments to multi-head attention models, such as GPT-2 (Radford et al., 2019), and verify that the insights drawn from the single-head model remain valid for more complex architectures.
@@ 2 Related Work
Challenges of transformers in compositional tasks. Recent advancements in large language models (LLMs) have showcased remarkable capabilities, often surpassing human performance (Fu et al., 2022; Wei et al., 2022). However, despite their impressive performance on single-step reasoning tasks (Srivastava et al., 2022), transformers struggle with multi-step compositional tasks and out-of-distribution (OOD) generalization (Csordás et al., 2021, 2022; Dziri et al., 2024; Hupkes et al., 2018; Lepori et al., 2023; Okawa et al., 2023; Yun et al., 2022). Ramesh et al. (Ramesh et al., 2023) show that transformers trained to directly compose capabilities struggle to generalize to OOD tasks with a synthetic task. Liu et al. (Liu et al., 2022) suggest that shallow transformers learn shortcuts during training, leading to poor OOD generalization. Numerous studies have explored various approaches to address these limitations, such as encouraging explicit reasoning step generation within a single generation (Wei et al., 2022), leveraging LLMs to generate reasoning steps iteratively (Creswell et al., 2022; Creswell and Shanahan, 2022). Despite these efforts, achieving complete mastery in compositional tasks remains a significant challenge for vanilla transformers. In order to clearly study the behaviors and internal mechanisms of language models, Zhang et al. Zhang et al. (2024) introduced anchor functions as benchmark functions for investigating transformer behavior. Our work builds on the anchor function setting to explore how different initialization scales affect model solutions and mechanisms.
@@
Parameter initialization of neural networks. The parameter initialization of the network is important to determine the final fitting result of the network (Arora et al., 2019; Chizat and Bach, 2018; E et al., 2020; Jacot et al., 2018; Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2018; Sirignano and Spiliopoulos, 2020; Williams et al., 2019; Zhang et al., 2019). Luo et al. (Luo et al., 2021), Zhou et al. (Zhou et al., 2022) mainly identify the linear regime and the condensed regime for two-layer and three-layer wide ReLU NNs, respectively. A series of works suggests that the condensed networks are often accompanied by good generalization ability of the model (Zhang et al., 2021, 2022, 2023; Zhang and Xu, 2024). Recent studies have also investigated the impact of initialization on the training of LLMs (Huang et al., 2020; Liu et al., 2020; Trockman and Kolter, 2023; Wang et al., 2024; Zhang et al., 2019; Zhu et al., 2021). These works primarily focus on how the initialization scale affects the stability of the training process and plays a crucial role in ensuring smooth and effective training of LLMs. In our work, we find that different initialization scales can significantly influence a model’s ability to memorize and infer data on the compositional task, highlighting the profound impact of initialization on the final performance and underlying mechanisms of the trained models.
@@ 3.5 Model Architecture and Basic Experimental Setups
$X^{\mathrm{ao}(l)}=\text{LN}(X^{(l)}+X^{\mathrm{qkv}(l)}W^{\mathrm{attn},l}),\quad X^{(l+1)}:=X^{\mathrm{do}(l)}=\text{LN}(\text{FNN}(X^{\mathrm{ao}(l)})+X^{\mathrm{ao}(l)}).$ (6)
The final output is obtained by projecting the output of the last layer $X^{\mathrm{do}(L)}$ using a linear projection layer, followed by a softmax operation and argmax to obtain the predicted token.
@@ 4 Two Phases of Solutions for Composite Functions
Figure 2: (a,b) Phase diagram of generalization performance on the unseen anchor (4, 3). (a) The model’s test accuracy based on the symmetric mapping. (b) The model’s test accuracy based on the inferential mapping. The abscissa represents the initialization rate $\gamma$, which corresponds to the standard deviation $(1/d_{\mathrm{in}})^{\gamma}$ of a normal distribution with a mean of 0 used for parameter initialization. The ordinate represents the depth of the transformer model. The shadow zones indicate the test accuracy on seen anchors is less than 90%. (c) Comparison of accuracy on the unseen anchor (4, 3) for both the inferential and symmetric solutions across different initialization rates $\gamma$ on GPT-2-SMALL. The error bars represent the standard deviation across 4-time runs.
@@
To validate the generality of these findings, we extend our experiments to the multi-head GPT-2-SMALL model. As shown in Fig. 2(c), the model exhibits similar generalization performance trends on the unseen pair (4, 3) across different initialization rates $\gamma$. The range of $\gamma$ variation is smaller in this setup, likely due to the difference in model parameter scale between the GPT-2 and the transformer models used in the previous experiments.
@@ 5.2 Divergence in Fused Vector Representations across Two Phases
As illustrated in Fig. 4(a, b), the heatmaps display the cosine similarity between output vectors for various anchor-key item pairs. Both the abscissa and ordinate represent key item values ranging from 30 to 40. The lower and upper triangles of the heatmap correspond to the cosine similarity matrices between $\bm{v}(x_{1};3,3)$ and $\bm{v}(x_{2};2,2)$, and between $\bm{v}(x_{1};1,2)$ and $\bm{v}(x_{2};1,3)$, respectively, where $x_{1}$ and $x_{2}$ are the corresponding key item values on the axes. The red boxes highlight positions where the inferential targets are the same, for example, the red highlight in the first column in Fig. 4(a) indicates $f(36;3,3)=f(30;2,2)$, where $f(\cdot;\cdot,\cdot)$ is defined in Equation 1.
@@ 5.3 Model Complexity: A Key Factor in Phase Transitions
Large initialization scales endow models with high complexity, allowing them to fit training data with minor parameter changes, as seen in FNNs in the linear regime (e.g., Neural tangent kernel) (Jacot et al., 2018; Luo et al., 2021). Conversely, models with small initialization scales start with low complexity and gradually increase it during training. In FNNs with small initialization, neurons in the same layer first condense in a few directions (Zhou et al., 2021), i.e., their input weights align and concentrate in a few directions, followed by condensing in more directions (Luo et al., 2021; Zhang et al., 2022, 2021), regularizing FNNs to fit data with minimal condensation directions. This phenomenon of parameter condensation is closely related to the model’s complexity and its ability to learn the underlying structure of the data. To better understand the mechanisms behind inferential and symmetric solutions, we investigate the degree of parameter condensation.
Figure 5: (a, b) Cosine similarity of neurons in the $W^{Q(1)}$ matrix. (a) Inferential solution with small initialization. (b) Symmetric solution with large initialization. (c, d) Visualization of the embedding space using T-SNE for different initialization scales. (c) Inferential solution with small initialization. The embedded tokens seem to form arithmetic sequences with common differences of 3 (red arrow) and 4 (blue arrow) along the two directions. (d) Symmetric solution with large initialization. Please refer to Appendix D for more detailed experimental results under different model depths and initialization rates $\gamma$.
@@
6 Prediction of mechanisms underlying different solutions
To measure the training steps required to fit data, we use the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. We conduct 9 random trials for each setting. As shown in Fig. 6, models that tend to learn data with inferential solutions indeed use more training steps for data with larger inferential complexity (Fig. 6(a)), while models that tend to learn data with symmetric solutions use roughly the same training steps for data with different inferential complexity (Fig. 6(b)). This prediction convincingly backs up our analysis of why different initialization scales can lead to different solutions. We further demonstrate the learning speed of each anchor pair in Appendix F.
@@ 7 Discussion
Conclusion. In this work, we investigate the influence of parameter initialization scale on the solutions learned by transformers for compositional tasks. We discover a phase diagram of model solutions, where small initialization scales lead to inferential solutions that capture the underlying compositional primitives, while larger initialization scales result in symmetric solutions that memorize mappings. To explain these findings, we analyze the information flow and vector representations within the model, revealing distinct mechanisms for inferential and symmetric solutions. Inferential solutions exhibit low complexity and learn individual mappings for single anchors, while symmetric solutions directly combine anchor information without learning the compositional structure. Building upon these insights, we can predict and experimentally confirm the differing effects of inferential complexity on the training speed of models with different initialization scales.
Limitation and Future Work. The key limitation of our work is that the experiments and analyses are based on synthetic data, which may not fully capture the complexities of real-world datasets and tasks. Although some of our conclusions have also been validated in models like GPT-2, further verification on a wider range of data and models is necessary. Going forward, we plan to extend our investigation to real-world datasets and tasks, to bridge the gap between our theoretical understanding and practical application. This could involve leveraging Mixture of Experts (MoE) to design networks with different initialization scales for different expert models, and exploring improved training strategies that adapt the model’s focus on enhancing memorization or inference capabilities at different stages of training with different initialization scales.

### v3

@@ Abstract
Transformers have shown impressive capabilities across various tasks, but their performance on compositional problems remains a topic of debate. In this work, we investigate the mechanisms of how transformers behave on unseen compositional tasks. We discover that the parameter initialization scale plays a critical role in determining whether the model learns inferential solutions, which capture the underlying compositional primitives, or symmetric solutions, which simply memorize mappings without understanding the compositional structure. By analyzing the information flow and vector representations within the model, we reveal the distinct mechanisms underlying these solution types. We further find that inferential solutions exhibit low complexity bias, which we hypothesize is a key factor enabling them to learn individual mappings for single anchors. We validate our conclusions on various real-world datasets. Our findings provide valuable insights into the role of initialization scale in shaping the type of solution learned by transformers and their ability to learn and generalize compositional tasks.
@@ 1 Introduction
Large-scale transformers (Vaswani et al., 2017) demonstrate unprecedented capabilities (Achiam et al., 2023; Brown et al., 2020; Choi et al., 2021; Teubner et al., 2023), even noted as “sparks of AGI" (Bubeck et al., 2023). However, their performance on compositional tasks, which are considered a key property of human cognition (Marcus, 2003; Smolensky et al., 2022), has long been the subject of intense debate, especially in terms of the out-of-distribution (OOD) generalization ability. This raises critical open questions about how to faithfully interpret transformers’ capabilities on compositional tasks. Can transformers learn the underlying compositional primitives within the data, or do they merely learn input-output mappings? When transformers produce incorrect outputs on compositional tasks, do their responses follow any discernible patterns or logic?
@@
Our work highlights the crucial role of initialization scale in shaping the solutions learned by transformers on compositional tasks, enabling them to better capture compositional structures and generalize to unseen tasks. Furthermore, adapting the initialization scale to the specific task type could be promising: larger scales for memorization tasks such as text memorization, and smaller scales for inference tasks such as code generation. Unless otherwise specified, our main text primarily focuses on a single-head attention model to facilitate a clearer understanding of the underlying mechanisms. We further extend our experiments to GPT-2 (Radford et al., 2019), and verify that the insights drawn from the single-head model remain valid for more complex architectures on various real-world datasets.
@@ 2 Related Work
Challenges of transformers in compositional tasks. Recent advancements in large language models (LLMs) have showcased remarkable capabilities, often surpassing human performance (Fu et al., 2022; Wei et al., 2022). However, despite their impressive performance on single-step reasoning tasks (Srivastava et al., 2022), transformers struggle with multi-step compositional tasks and OOD generalization (Csordás et al., 2021, 2022; Dziri et al., 2024; Hupkes et al., 2018; Lepori et al., 2023; Okawa et al., 2023; Yun et al., 2022). Ramesh et al. (Ramesh et al., 2023) show that transformers trained to directly compose capabilities struggle to generalize to OOD tasks with a synthetic task. Liu et al. (Liu et al., 2022) suggest that shallow transformers learn shortcuts during training, leading to poor OOD generalization. Numerous studies have explored various approaches to address these limitations, such as encouraging explicit reasoning step generation within a single generation (Wei et al., 2022), leveraging LLMs to generate reasoning steps iteratively (Creswell et al., 2022; Creswell and Shanahan, 2022). Despite these efforts, achieving complete mastery in compositional tasks remains a significant challenge for vanilla transformers. A series of works studies the internal mechanisms of language models and improves the capabilities of language models (Wang et al., 2024a, b, 2023; Cao et al., 2024). In order to clearly study the behaviors and internal mechanisms of language models, Zhang et al. Zhang et al. (2024) introduced anchor functions as benchmark functions for investigating transformer behavior. Our work builds on the anchor function setting to explore how different initialization scales affect model solutions and mechanisms.
@@
Parameter initialization of neural networks. The parameter initialization of the network is important to determine the final fitting result of the network (Arora et al., 2019; Chizat and Bach, 2018; Zhang et al., 2019; E et al., 2020; Jacot et al., 2018; Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2018; Sirignano and Spiliopoulos, 2020; Williams et al., 2019). Luo et al. (Luo et al., 2021), Zhou et al. (Zhou et al., 2022) mainly identify the linear regime and the condensed regime for two-layer and three-layer wide ReLU NNs, respectively. A series of works suggests that the condensed networks are often accompanied by good generalization ability of the model (Zhang et al., 2022; Zhang and Xu, 2023; Zhang et al., 2023; Zhang and Xu, 2024). A line of works links the grokking phenomenon with the improvement of reasoning ability (Power et al., 2022; Wang et al., 2024; Gopalani et al., 2024) and points out that the initialization scale has an important influence on the occurrence of grokking (Liu et al., 2022). Recent studies also investigate the impact of initialization on the training process of LLMs (Huang et al., 2020; Liu et al., 2020; Trockman and Kolter, 2023; Wang et al., 2024; Zhang et al., 2019; Zhu et al., 2021). These works primarily focus on how the initialization scale affects the stability of the training process and plays a crucial role in ensuring smooth and effective training of LLMs. In our work, we find that different initialization scales can significantly influence a model’s ability to memorize and infer data on the compositional task, highlighting the profound impact of initialization on the final performance and underlying mechanisms of the trained models.
@@ 3.5 Model Architecture and Basic Experimental Setups
$X^{\mathrm{ao}(l)}=\text{LN}(X^{(l)}+X^{\mathrm{qkv}(l)}W^{\mathrm{attn},l}),\quad X^{(l+1)}:=X^{\mathrm{do}(l)}=\text{LN}(\text{FNN}(X^{\mathrm{ao}(l)})+X^{\mathrm{ao}(l)}),$ (6)
where “LN” refers to Layer Normalization and “FN” represents a Fully Connected Network. The final output is obtained by projecting the output of the last layer $X^{\mathrm{do}(L)}$ using a linear projection layer, followed by a softmax operation and argmax to obtain the predicted token.
@@ 4 Two Phases of Solutions for Composite Functions
Figure 2: (a,b) Phase diagram of generalization performance on the unseen anchor (4, 3). (a) The model’s test accuracy based on the symmetric mapping. (b) The model’s test accuracy based on the inferential mapping. The abscissa represents the initialization rate $\gamma$, which corresponds to the standard deviation $(1/d_{\mathrm{in}})^{\gamma}$ of a normal distribution with a mean of 0 used for parameter initialization. The ordinate represents the depth of the transformer model. The shadow zones indicate the test accuracy on seen anchors is less than 90%. (c) Comparison of accuracy on the unseen anchor (4, 3) for both the inferential and symmetric solutions across different initialization rates $\gamma$ on GPT-2. The error bars represent the standard deviation across 4-time runs.
@@
To validate the generality of these findings, we extend our experiments to the multi-head GPT-2 model. As shown in Fig. 2(c), the model exhibits similar generalization performance trends on the unseen pair (4, 3) across different initialization rates $\gamma$. The range of $\gamma$ variation is smaller in this setup, likely due to the difference in model parameter scale between the GPT-2 and the transformer models used in the previous experiments.
@@ 5.2 Divergence in Fused Vector Representations across Two Phases
As illustrated in Fig. 4(a, b), the heatmaps display the cosine similarity between output vectors for various anchor-key item pairs. Both the abscissa and ordinate represent key item values ranging from 30 to 40. The lower and upper triangles of the heatmap correspond to the cosine similarity matrices between $\bm{v}(x_{1};3,3)$ and $\bm{v}(x_{2};2,2)$, and between $\bm{v}(x_{1};1,2)$ and $\bm{v}(x_{2};1,3)$, respectively, where $x_{1}$ and $x_{2}$ are the corresponding key item values on the axes. The red boxes highlight positions where the inferential targets are the same, for example, the red highlight in the first column in Fig. 4(a) indicates $f(36;3,3)=f(30;2,2)$, where $f(\cdot;\cdot,\cdot)$ is defined in Equation (1).
@@ 5.3 Model Complexity: A Key Factor in Phase Transitions
Large initialization scales endow models with high complexity, allowing them to fit training data with minor parameter changes, as seen in FNNs in the linear regime (e.g., Neural tangent kernel) (Jacot et al., 2018; Luo et al., 2021). Conversely, models with small initialization scales start with low complexity and gradually increase it during training. In small initialization FNNs, the input weights of different neurons often cluster along a few isolated orientations, which is referred to as condensations (Zhou et al., 2021; Zhang et al., 2021, 2022). This phenomenon of parameter condensation is closely related to the model’s complexity and its ability to learn the underlying structure of the data. To better understand the mechanisms behind inferential and symmetric solutions, we investigate the degree of parameter condensation.
Figure 5: (a, b) Cosine similarity of neurons in the $W^{Q(1)}$ matrix. The abscissa and ordinate both represent neuron index. (a) Inferential solution with small initialization. (b) Symmetric solution with large initialization. (c, d) Visualization of the embedding space using t-SNE for different initialization scales. (c) Inferential solution with small initialization. The embedded tokens seem to form arithmetic sequences with common differences of 3 (red arrow) and 4 (blue arrow) along the two directions. (d) Symmetric solution with large initialization. Please refer to Appendix D for more detailed experimental results under different model depths and initialization rates $\gamma$.
@@
6 Further Verification on Realistic Tasks
We validated the performance of models with different initialization scales and weight decay settings across a series of compositional and reasoning tasks. Below, we introduce each task and the corresponding results. Please refer to the Appendix F for a detailed introduction of each dataset.
Compositional tasks: SCAN and COGS. For the SCAN dataset (Lake and Baroni, 2018), we selected the “Generalizing composition across primitive commands" task, where the “turn left" command only appears in single-command mappings and is trained alongside other composite commands. We assess the model’s generalization ability on composite commands that include the “turn left" command. For the COGS dataset (Kim and Linzen, 2020), we evaluate the model’s in-distribution and out-of-distribution generalization performance after training on the same training set.
Figure 6: Performance comparison of models with different initialization scales and weight decay coefficients on compositional tasks. (a) For the SCAN task, we assess the generalization ability on composite commands that include the “turn left" command. (b, c) For the COGS task, we evaluate (b) in-distribution and (c) out-of-distribution generalization after training on the same dataset. Small initialization and large weight decay (blue) consistently outperform large initialization and small weight decay (orange) across different tasks and data scales. The parameters are initialized following a zero-mean normal distribution with a standard deviation of $d_{in}^{-\gamma}$.
As shown in Fig. 6, we display the generalization performance of models with different initialization scales and weight decay coefficients across various data sizes. Small initialization and large weight decay consistently outperform large initialization and small weight decay across different task types and data scales. Notably, in the COGS task, even when the in-distribution generalization of both settings (with 20k training data) reaches over 99%, the difference in out-of-distribution generalization remains significant.
Reasoning tasks: PrOntoQA. PrOntoQA (Saparov and He, 2023) is a synthetic multi-step reasoning dataset where each data point assigns hierarchical relationships among objects and requires the model to determine whether a multi-step reasoning chain is correct. Fig. 7 illustrates the convergence rates and generalization errors [^2] During testing, we only evaluate the model’s accuracy in judging hierarchical relationships. Thus, the model’s random guessing accuracy is 50%. with respect to data scale for models with large initialization (and small weight decay, Fig. 7(a)) and small initialization (and large weight decay, Fig. 7(b)). An interesting phenomenon is observed for models with large initialization (small weight decay): as the data size increases, the convergence rate first decreases and then increases. When the data size is small, the model tends to fit the data through memorization. Therefore, as the data size increases, the training difficulty increases (i.e., the training speed slows down), and the model’s generalization ability is poor. As the data size grows further, the model, constrained by its complexity, can no longer memorize all the data and thus shifts to fitting the data through reasoning. This leads to an increase in fitting speed and results in better generalization.
Figure 7: Performance comparison of models with different initialization scales and weight decay coefficients on PrOntoQA. (a) Convergence steps and test accuracy for large initialization ($\gamma=0.5$) and small weight decay (WD = 0.01). (b) Convergence steps and test accuracy for small initialization ($\gamma=0.7$) and large weight decay (WD = 0.1). Models with large initialization initially struggle with memorization before improving as data size increases, whereas small initialization models maintain faster convergence and better generalization. The parameters are initialized following a zero-mean normal distribution with a standard deviation of $d_{in}^{-\gamma}$.
In contrast, models with small initialization (large weight decay) inherently prefer to fit the data through reasoning, leading to faster convergence and better generalization at the same data scale compared to models with large initialization (small weight decay).
@@ 7 Discussion
Conclusion. In this work, we investigate the influence of parameter initialization scale on the solutions learned by transformers for compositional tasks. We discover a phase diagram of model solutions, where small initialization scales lead to inferential solutions that capture the underlying compositional primitives, while larger initialization scales result in symmetric solutions that memorize mappings. To explain these findings, we analyze the information flow and vector representations within the model, revealing distinct mechanisms for inferential and symmetric solutions. Inferential solutions exhibit low complexity and learn individual mappings for single anchors, while symmetric solutions directly combine anchor information without learning the compositional structure. We further extend our experiments to GPT-2, and verify that the insights remain valid for more complex architectures on various real-world datasets.
Limitation and Future Work. The key limitation of our work is that the experiments and analyses are based on synthetic data, which may not fully capture the complexities of real-world datasets and tasks. Although some of our conclusions have also been validated in models like GPT-2, further verification on a wider range of models is necessary. Going forward, we plan to extend our investigation to real-world datasets and tasks, to bridge the gap between our theoretical understanding and practical application. This could involve leveraging Mixture of Experts (MoEs) to design networks with different initialization scales for different expert models, and exploring improved training strategies that adapt the model’s focus on enhancing memorization or inference capabilities at different stages of training with different initialization scales.
Acknowledgments
This work is sponsored by the National Key R&D Program of China Grant No. 2022YFA1008200, the National Natural Science Foundation of China Grant No. 92270001(Z. X.), 12371511 (Z. X.), 12422119 (Z. X.), 12101402 (Y. Z.), the Lingang Laboratory Grant No. LG-QS-202202-08 (Y. Z.), Shanghai Municipal of Science and Technology Major Project No. 2021SHZDZX0102 (Z. X., Y. Z.), and the HPC of School of Mathematical Sciences and the Student Innovation Center, and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University, Key Laboratory of Marine Intelligent Equipment and System, Ministry of Education, P.R. China. This work was partially supported by SJTU Kunpeng&Ascend Center of Excellence.